OpenStudy (anonymous):

Use implicit differentiation to find the slope of the tangent line to the curve 4x^2+4xy+3y^2=27 at the point (2,1) m=?

OpenStudy (amistre64):

power, product, power, constant rules

OpenStudy (amistre64):

treat each variable as a function in its own right

OpenStudy (anonymous):

I got this so far not sure if it's right y'(x)= -x-2y/2x+3y

OpenStudy (amistre64):

4x^2+4xy+3y^2=27 8xx' + 4x'y + 4xy' + 6yy' = 0 ; x'=1 8x + 4y + y'(4x + 6y) = 0 y' = -8x+4y ------- 4x+6y

OpenStudy (amistre64):

i wouldnt worry bout simplifyiing; just insert the x and y and then work the numbers

OpenStudy (amistre64):

and i missed a - :)

OpenStudy (amistre64):

maybe a -4x on yours would be better since you simplified; 8x halves to 4x

OpenStudy (anonymous):

Oh ok thanks for the help!

OpenStudy (amistre64):

yw

OpenStudy (anonymous):

oh what do you mean when u said 'a' ? sorry

OpenStudy (amistre64):

i meant that i missed this negative; had it as a + v y' = -8x-4y ------- 4x+6y

OpenStudy (anonymous):

Oh ok I see thanks Manh u really help me!